OpenStudy (anonymous):

Can somebody help me solve this? D: 5x + 10 = x + 50

OpenStudy (anonymous):

I don't know if I'm doing this right, because I always end up with a decimal.. o.O

OpenStudy (anonymous):

combine like terms first. :)

OpenStudy (anonymous):

I did all of that. but I always end up with a decimal.. It has to be a whole number. Dx

OpenStudy (anonymous):

Subtract 10 on Both sides to get you started

OpenStudy (anonymous):

okay so show me what u did :)

OpenStudy (anonymous):

5x + 10 = x + 50 10 = 6x + 50 -40 = 6x -6.66 = x I probably did this wrong, lol.

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

@sonymelo can you show me how you got to 10?

OpenStudy (anonymous):

subtract x from both sides subtract 10 from both sides

OpenStudy (anonymous):

this would have been an easier way to do it...

OpenStudy (anonymous):

5x + 10 = x + 50 That means five times a number plus ten is the same as one time that number plus fifty. 5 times 10 = 50, plus 10 = 60 10 + 50 = 60 60 = 60 x = 10.

OpenStudy (anonymous):

so you don't have all those negatives.

OpenStudy (anonymous):

5x + 10= x + 50 -x -x 4x + 10 = 50 -10 -10 4x = 40 x= 10

OpenStudy (anonymous):

..now I'm really confused. o.O

OpenStudy (anonymous):

u added an x where u should have subtracted

OpenStudy (anonymous):

 do you understand that?